### v1

# AI Agentic Vulnerability Injection And Transformation with Optimized Reasoning

Abstract
The increasing complexity of software systems and the sophistication of cyber-attacks have underscored the critical need for effective automated vulnerability detection and repair systems. Traditional methods, such as static program analysis, face significant challenges related to scalability, adaptability, and high false-positive and false-negative rates. AI-driven approaches, particularly those using machine learning and deep learning models, show promise but are heavily reliant on the quality and quantity of training data. This paper introduces a novel framework designed to automatically introduce realistic, category-specific vulnerabilities into secure C/C++ codebases to generate datasets. The proposed approach coordinates multiple AI agents that simulate expert reasoning, along with function agents and traditional code analysis tools. It leverages Retrieval-Augmented Generation for contextual grounding and employs Low-Rank approximation of weights for efficient model fine-tuning. Our experimental study on 116 code samples from three different benchmarks suggests that our approach outperforms other techniques with regard to dataset accuracy, achieving between 89% and 95% success rates in injecting vulnerabilities at function level.
Index Terms:
Automated Vulnerability Detection, AI-Driven Security, Agentic Workflows
I Introduction
Software vulnerabilities refer to weaknesses in system security requirements, design, implementation, or operation, that could be accidentally triggered or intentionally exploited, resulting in a violation of the system’s security policy [1]. The rapid growth in software complexity, coupled with the sophistication of cyber-attacks, poses a significant threat to the global security and stability of digital infrastructures. In 2024 alone, the total number of publicly reported vulnerabilities increased by 25% [2].
Consequently, effective automated vulnerability detection and repair systems have become critically important. Traditional methods, such as static program analysis, face significant challenges related to scalability, adaptability, and high false-positive and false-negative rates [3] [4]. Due to their reliance on manually crafted rule-based heuristics and the combinatorial complexity of their analysis, these programmatic approaches inherently constrain opportunities for potential significant breakthroughs.
AI-driven approaches for code vulnerability detection [5, 6, 7, 8, 9, 10, 11] and repair [12, 13, 14] have recently attracted considerable attention in the research community. Data-driven methods, especially those using machine learning and deep learning models, show promise in identifying and addressing a wide range of vulnerabilities. However, like most AI-driven systems, their effectiveness heavily relies on the quality and the quantity of the training data [11]. This dependency has become even more critical with the emergence of Large Language Models (LLMs), which can have billions of parameters to train. Nevertheless, the scarcity of large, high-quality training datasets has become a major obstacle to the progress of these techniques [11] [6] [5]. Existing vulnerability datasets can be grouped into the following categories based on how the code samples were collected and annotated:
  • Synthetic datasets: Both the code and its vulnerability labels are generated artificially, typically using predefined insecure coding patterns. A notable example is the Juliet test suite [15] which is a collection of small synthetic test cases, each demonstrating a specific weakness. It is part of a larger SARD dataset [16] which contains test cases ranging from small synthetic code to production software.
    These datasets offer clean and controllable examples but often lack the structural and semantic complexity of real-world code. As a result, AI models trained on such data overfit to superficial patterns—such as specific variable names or function structures—rather than learning the underlying causes of vulnerabilities [5].
  • Large realistic datasets: Several widely used large-scale vulnerability datasets, including ReVeal [5], BigVul [17], CVEfixes [18], CrossVul [19] and Diversevul [10] rely on heuristic-based labeling functions. These approaches typically involve identifying vulnerability-fixing commits from public security databases, such as the National Vulnerability Database (NVD), and labeling the pre-commit versions of affected functions as vulnerable. While these datasets contain code samples from real-world production software, the reliance on heuristics introduces noise and label inaccuracies. Studies have reported label accuracies as low as 25% for some datasets [11] [20]. Additionally, datasets like D2A [21] and Draper [22] use static analyzers to label code, which results in numerous false positives due to the limitations of static analysis tools. The presence of noisy or incorrect labels in all of these datasets poses challenges for training reliable vulnerability detection models [10] [11].
  • Reliable labels datasets: To enhance label accuracy, some datasets employ manual expert annotation or stricter heuristics. For instance, SVEN [23] manually labels 1,606 functions, improving label quality but limiting scalability due to resource constraints and coverage of only 9 types of Common Weakness Enumeration (CWE). Alternatively, PrimeVul [11] applied rigorous heuristics, claiming human-level labeling accuracy across  7,000 vulnerable and  229,000 benign functions spanning over 140 CWEs. Despite these advancements, the creation of larger high-quality datasets remains challenging to scale, especially when corresponding pairs of vulnerable and benign code samples are required.
  • Vulnerability injection-based datasets: Several approaches for injecting bugs into production software have been proposed and implemented [24, 25, 26, 27]. Methods such as VULGEN [28] and VinJ [29] combine pattern mining with deep learning-based localization to identify injection sites and apply vulnerability-inducing edits. These approaches scale well and produce diverse, realistic samples. However, injected bugs may lack the semantic and contextual richness of real-world vulnerabilities, and current techniques typically support only a limited set of CWE types. Moreover, the accuracy of injected datasets remains a concern: for example, VULGEN achieves only a 69% success rate in generating truly vulnerable samples, indicating potential noise and limited reliability.
To address these challenges, we introduce AVIATOR (AI Agentic Vulnerability Injection And Transformation with Optimized Reasoning), an AI agentic workflow designed to automatically introduce realistic, category-specific vulnerabilities into secure C/C++ codebases. AVIATOR coordinates multiple AI agents that simulate expert reasoning, along with function agents and traditional code analysis tools. It leverages Retrieval-Augmented Generation (RAG) [30] for contextual grounding and employs Low-Rank Adaptation (LoRA) [31] for efficient model fine-tuning.
To evaluate the performance of AVIATOR in introducing vulnerabilities we conducted an experimental study on 116 code samples from three different benchmarks used in the literature [16, 32, 11]. We leveraged formal verification tool ESBMC [33] for supported vulnerabilities and manual verification by experienced computer security researchers. We compare AVIATOR with existing automated large-scale realistic software vulnerability dataset creation technique. The results suggest a potential for the AVIATOR to outperform other techniques with regard to accuracy of dataset labels. It achieved between 89% and 95% success rates in injecting vulnerabilities across the different benchmarks.
Our contributions can be summarized as follows:
  • Agentic Injection Workflow: To the best of our knowledge, we are the first to develop a modular, multi-agent pipeline that replicates the reasoning processes of cybersecurity experts to generate plausible vulnerability injections.
  • Contextual Augmentation: By integrating Retrieval-Augmented Generation (RAG) modules, we ensure that vulnerability transformations are grounded in diverse, real-world code contexts.
  • Efficient Fine-Tuning: We utilize Low-Rank Adaptation (LoRA) to fine-tune large language models on small-scale datasets with minimal computational cost, thereby enhancing scalability.
  • Empirical Validation: Our preliminary evaluation demonstrates that AVIATOR achieves between 89% and 95% success in generating plausible vulnerabilities, facilitating the creation of large-scale, accurately labeled, and diverse datasets for training advanced detection models. Through our experiments, we have established the defining characteristics of high-quality software vulnerability datasets, essential for both training AI models and benchmarking purposes.
The remainder of this paper is organized as follows: Section 2 reviews existing methods for vulnerability detection and data generation. Section 3 details the architecture and agentic workflow of AVIATOR. Section 4 outlines our experimental setup and evaluation metrics. Section 5 presents empirical results and analysis. Section 6 discusses limitations and future work, and Section 7 concludes.
II Background and Motivation
In this section, we introduce the fundamental concepts and challenges associated with software vulnerability injection to motivate our proposed approach. We first describe the properties necessary for constructing effective datasets to train and benchmark AI-based vulnerability detection and repair systems. Then, we formalize and illustrate the process of vulnerability injection in source code. Additionally, we clarify the scope and objectives that underpin our proposed solution.
II-A Properties of High-Quality Vulnerability Datasets
Effective training and rigorous benchmarking of AI-based vulnerability detection and repair methods critically depend on the quality of the underlying datasets. To clearly define the desired dataset characteristics, we introduce the ABCD criteria, similar to [1] and [25]:
  • Accurate: Labels must reflect ground truth with clear information on the type of vulnerability, specifically CWE identifiers and, when possible, precise vulnerability locations—essential for reliable model training and benchmark evaluation.
  • Big: The dataset should be large enough to support training of deep models and enable statistically meaningful benchmarking of detection and repair systems.
  • Credible: Code samples should reflect the structure and complexity of real-world software to ensure model generalization in practical settings.
  • Diverse: The dataset should cover varied coding styles and vulnerability types to enable generalizable learning and comprehensive benchmarking across scenarios.
In practice, additional constraints improve training utility:
  • Each sample should contain a single vulnerability type to avoid ambiguity during training and evaluation.
  • Paired samples, benign and vulnerable versions, are ideal for repair tasks and contrastive learning.
II-B Motivation
Creating datasets that rigorously fulfill the ABCD criteria and structural requirements remains an open challenge due to significant limitations inherent to existing approaches [6]. Synthetic datasets, while accurate and easily scalable, lack credibility and offer limited diversity, as they rely on simplistic, predefined vulnerability patterns [5]. Heuristic-based labeling, in contrast, provides credible and diverse samples by mining real-world open-source repositories, but it suffers from a trade-off between accuracy and scalability: stricter heuristics yield smaller but cleaner datasets [11], while relaxed heuristics introduce noise [10]. Static analysis-based labeling can generate large datasets, yet accuracy remains a concern due to high false-positive rates. Manual labeling, though highly accurate, is not scalable for large datasets [34].
Vulnerability injection presents a promising alternative. By design, it enables scalability through automation and preserves credibility when applied to real-world codebases. However, the effectiveness of this approach depends heavily on the injection technique’s ability to introduce diverse and accurate vulnerabilities. Current state-of-the-art methods remain limited in this regard: recent studies report coverage of only 19 distinct CWE classes and achieve an injection success rate of just 69% [28, 29]. Such limitations constrain both the diversity and reliability of resulting datasets, particularly for benchmarking purposes. Nonetheless, these initial results suggest that with more refined techniques, vulnerability injection has the potential to systematically produce accurate, realistic, diverse, and large-scale datasets that better support the development and test of modern AI-based vulnerability detection and repair systems, as well as traditional static analyzers.
II-C Vulnerability Injection
We consider a code snippet $c$ as an ordered sequence of lexical tokens $c=\langle c_{1},c_{2},\dots,c_{n}\rangle,\quad c_{i}\in T$, where $T$ is the vocabulary of the programming language (e.g., operators, delimiters, identifiers, literals, keywords, punctuation), and the token sequence conforms to the language grammar.
We define a vulnerability injection operator:
$\mathsf{inject}:T^{n}\times V\longrightarrow T^{m},\qquad\bigl{(}c^{b},v_{j}\bigr{)}\mapsto c^{\mathrm{v}}$ (1)
which transforms a benign snippet $c^{b}$ into a vulnerable snippet $c^{v}$ exhibiting the vulnerability $v_{l}\in V$, where $V$ is the set of existing vulnerability types.
Injection is performed through a finite set of token-level edits. Let $I=\{\,i_{1},\dots,i_{k}\}\subseteq\{1,\dots,n\}$ be the set of positions to modify, and let each edit be defined by a function
$\tau_{i}:T\longrightarrow T^{*},\quad\tau_{i}(c^{b}_{(i)})=\langle t^{(i)}_{1},\dots,t^{(i)}_{m^{(i)}}\rangle,$
where $m^{(i)}\geq 0$, and $T^{*}$ is the set of finite token sequences. This formulation captures:
  • Insertion: New tokens are inserted before or after $c^{b}_{(i)}$:
[…]
  • Deletion: $c^{b}_{(i)}$ is removed entirely:
    $\quad\tau_{i}(c^{b}_{(i)})=\langle\rangle,\quad m^{(i)}=0.$
  • Replacement: $c^{b}_{(i)}$ is replaced by a new non-empty token sequence:
    $\tau_{i}(c^{b}_{(i)})=\langle t^{(i)}_{1},\dots,t^{(i)}_{m^{(i}}\rangle,\quad m^{(i)}\geqslant 1.$
For positions $i\notin I$, we define $\tau_{i}(c^{b}_{(i)})=\langle c^{b}_{(i)}\rangle$, leaving the token unchanged.
The final injected sequence is obtained by flattening all transformed segments:
$c^{v}=\bigoplus_{i=1}^{n}\tau_{i}(c^{b}_{(i)})=\langle c^{\prime}_{1},c^{\prime}_{2},\dots,c^{\prime}_{m}\rangle,$ (2)
where $\bigoplus$ denotes sequence concatenation.
A high-quality injection must satisfy the following properties:
  • Realism: Edits $\{\tau_{i}\}$ should reflect patterns observed in real-world vulnerabilities of class $v_{j}$ (e.g. adapt to existing dataflow, meaningful output, no dead storage …).
  • Functional Integrity: $c^{v}$ must remain syntactically and semantically valid, preserving the original functionality of $(c^{b})$, aside from the injected flaw.
  • Minimality: The number of edits $|I|$ should be kept minimal.
Figure 1: Architecture of the AVIATOR framework illustrating the two core modules. The top module represents the Vulnerability Injection component, which utilizes RAG and supervised fine-tuning with LoRA for guiding code transformations. The bottom module depicts the Validation component, employing LLM-based discriminators and static analysis to verify the presence and accuracy of injected vulnerabilities.
III Proposed Framework
Existing vulnerability injection techniques continue to face fundamental limitations in both coverage and reliability. These methods generally support only a limited range of vulnerability types and exhibit low success rates in injection. This is largely due to their inability to adapt to the structural and contextual intricacies of real-world code found in production, as well as a lack of deep semantic reasoning.
To overcome these limitations, the proposed AVIATOR framework comprises two core modules: (1) a vulnerability injection module that utilizes retrieval-augmented generation (RAG) and supervised fine-tuning with Low-Rank Adaptation (LoRA) to guide code transformations; and (2) a validation module that employs LLM-based discriminators and a static analyzer to meticulously verify the presence and accuracy of the injected vulnerabilities. These two modules are illustrated in Figure 1.
In the following, we present the AVIATOR framework’s core components and methodologies. We first define AI agentic workflows, where specialized agents collaborate through structured execution paths to enhance modularity and error correction, mimicking cybersecurity experts’ analytical processes. We then detail the Vulnerability Injection Module of the workflow, which transforms secure code into vulnerable code using semantic analysis and advanced techniques like retrieval-augmented generation and fine-tuning. Additionally, we discuss the Validation Module, which ensures the accuracy and presence of injected vulnerabilities through a combination of deterministic tools and language model agents.
III-A AI Agentic Workflows
Agentic workflows are systems in which a collection of agents are orchestrated to perform a complex task by following a predefined execution path [35]. Each agent is responsible for a specific action and transforms structured inputs into outputs that serve subsequent agents. Agents can be instantiated as large language models (LLMs) or deterministic programmatic tools (e.g., static analyzers).
Formally, we define an agentic workflow $W_{k}$ of $k$ agents as a directed execution graph, where each node represents an agent and edges denote conditional transitions. Each agent $A_{i}$ transforms structured inputs into outputs, given by $A_{i}:X_{i}\times E\mapsto Y_{i}$, with $E$ denoting external contextual artifacts (e.g., retrieved knowledge). Transitions depend on conditions evaluated dynamically from prior outputs.
This architectural paradigm offers advantages over monolithic reasoning or linear pipelines. First, it enables modular decomposition: the overall task is split into smaller, logically coherent subtasks aligned with the competencies of individual agents. This reduces the reasoning burden on any single component and improves interpretability. Second, it allows structured context propagation: each agent receives only the relevant inputs for its subtask, while outputs from earlier agents can be enriched with retrieval or validation results. Third, agentic workflows support corrective execution: if an agent detects inconsistencies or failures in prior outputs, the workflow can re-invoke earlier agents to revise decisions or recover from errors.
In AVIATOR, we implement an agentic workflow that mirrors the step-by-step reasoning of a cybersecurity analyst. The system orchestrates specialized agents responsible for tasks such as vulnerability selection, injection point identification, transformation synthesis, and post-injection validation, as detailed in III-B, III-C. This modular design promotes targeted decision-making, minimizes error propagation, and ensures that vulnerability injection proceeds in a systematic, verifiable, and context-aware manner.
III-B Vulnerability Injection Module
The Vulnerability Injection Module in AVIATOR plays a crucial role by converting a benign code snippet into a version that includes a specific vulnerability. Central to the framework, this module ensures that the injected vulnerability is not only present but also realistic, minimal, and contextually coherent. It comprises the following components.
III-B1 Semantic Analysis and Injection Point Identification
The process begins with semantic and structural analysis of the benign code. An LLM-based agent analyzes the function’s control and data flow to infer its purpose, input/output behavior, and internal logic (Code Understanding agent). It then qualitatively identifies potential sources—locations where untrusted input may enter the program—and sinks—locations where sensitive operations occur (Sanitization Detector agent). The agent also detects sanitization logic that would prevent the intended vulnerability from manifesting, such as bounds checks or input validation.
This process mirrors traditional symbolic static analysis and taint-tracking techniques [36], but applied inversely: rather than proving the presence of a vulnerability, we infer the barriers that prevent it, and identify how they could be weakened or removed. This reverse reasoning is critical to selecting feasible and non-trivial injection points.
III-B2 RAG-Based Vulnerability Injection with Fine-Tuning
In the following, we first explain the retrieval-augmented vulnerability generation, then detail the low-rank adaptation fine-tuning method, followed by supervised fine-tuning and reinforcement learning strategies.
Retrieval-Augmented Vulnerability Generation
Once candidate injection points are identified, an agent transforms the code to inject the target vulnerability (Vulnerability Injector agent). This agent operates under a retrieval-augmented generation (RAG) setup. Given an input function $c^{b}$, the agent retrieves a set of similar benign/vulnerable pairs $\{c^{b_{(j)}},c^{v_{(j)}}\}_{j=1}^{k}$ from a knowledge base using embedding-based similarity. Retrieved examples are aligned using line-level diffs, marking the injected edits.
This retrieved context alongside the analyzed output, is provided to the LLM to guide code transformation. This ensures that the injection edits reflect realistic code structure and vulnerability patterns observed in real-world examples.
Low-Rank Adaptation (LoRA)
To specialize the vulnerability injection agent without incurring the computational cost of full model fine-tuning, we apply Low-Rank Adaptation (LoRA) [31]. This technique is particularly effective in settings where the base model is large, but the task-specific dataset is limited in size.
Rather than updating the full weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA learns a low-rank decomposition by injecting two trainable matrices $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$, with $r\ll\min(d,k)$, into the model’s architecture: $W=W_{0}+BA$.
This reduces the number of trainable parameters while retaining the benefits of large-scale pretraining.
We explore two distinct fine-tuning paradigms for the vulnerability injection agent: supervised fine-tuning (SFT) and reinforcement learning with Group Relative Policy Optimization (GRPO). Both approaches operate on the same dataset of aligned benign/vulnerable function pairs:
$\mathcal{Q}=\left\{\left(c^{b_{(j)}},\,\hat{c}^{v_{(j)}}\right)\right\}_{j=1}^{N}$ (3)
where each benign function $c^{b_{(j)}}$ is accompanied by a structured prompt $q^{(j)}$. The prompt is generated by prior workflow steps from III-B1, incorporating (1) semantic and structural analysis results of the benign code, (2) the original benign function itself, and (3) relevant examples from the Retrieval-Augmented Generation (RAG) knowledge base, annotated to highlight vulnerability injection locations. The objective is for the agent to generate the corresponding vulnerable function $\hat{c}^{v}_{j}$ given this prompt.
Supervised Fine-Tuning (SFT)
The vulnerability injection agent is first trained via supervised fine-tuning (SFT), using a standard autoregressive next-token prediction objective. Let $q^{(j)}$ denote the structured prompt. For example $c^{b_{(i)}}$ and $\hat{\mathbf{c}}^{{v}_{i}}=\langle y^{(i)}_{1},\dots,y^{(i)}_{T^{(i)}}\rangle$ the tokenized target vulnerable function of size $T^{(i)}$. We fine-tune the LoRA adapter parameters $\phi$ by minimizing the negative log-likelihood of each next token under teacher forcing. The loss for a batch of size $B$ is:
$\mathcal{L}_{\text{SFT}}(\phi)=-\frac{1}{B}\sum_{i=1}^{B}\sum_{t=0}^{T_{i}-1}\log P\bigl{(}y_{i,t+1}\,\bigm{|}\,\mathbf{x}_{i},\,y_{i,1},\dots,y_{i,t};\,\phi\bigr{)}.$ (4)
Here, $P(\cdot\mid\cdot;\phi)$ is the autoregressive distribution defined by the LoRA-adapted LLM. This technique focuses on exact token-level matching with the aim that it will learn underlying concepts behind vulnerability injection, guided explicitly by ground-truth examples.
Group Relative Policy Optimization (GRPO)
While SFT optimizes for token-level accuracy, it does not directly focus on structural and semantic nuances of code. Therefore, we also apply reinforcement learning (RL) via Group Relative Policy Optimization (GRPO) [37], using the CodeBLEU [38] metric as a reward function.
GRPO is an efficient reinforcement learning technique that eliminates the need for a separate critic network by comparing multiple candidate outputs for each prompt and computing their relative rewards within a sampled group. Given a prompt $q$, we sample $G$ candidate outputs $\{o_{1},o_{2},\dots,o_{G}\}$ from the old policy model $\pi_{\theta_{\text{old}}}$. Each candidate output $o_{i}$ receives a reward score $r(o_{i})$ based on the CodeBLEU metric, which better captures structural and semantic properties of code. CodeBLEU combines four components:
  • N-gram match: standard BLEU score over tokens.
  • Weighted n-gram match: gives higher importance to syntax-critical tokens.
  • AST match: measures similarity of abstract syntax trees.
  • Data-flow match: compares data-flow graphs to assess semantic equivalence.
$\text{CodeBLEU}=\alpha\cdot\text{BLEU}(n\text{-grams})+\beta\cdot\text{BLEU}_{w}(n\text{-grams})\\ +\gamma\cdot\text{Match}_{\text{AST}}+\delta\cdot\text{Match}_{\text{Data-flow}},$ (5)
with $\alpha+\beta+\gamma+\delta=1$. By default, $\alpha=\beta=\gamma=\delta$ [38].
The normalized advantage for each candidate output is computed as:
$A_{i}=\frac{r(o_{i})-\mu}{\sigma},\quad\mu=\frac{1}{G}\sum_{j=1}^{G}r(o_{j}),\quad\sigma=\text{std}\left(\{r(o_{j})\}_{j=1}^{G}\right).$ (6)
GRPO optimizes the policy model $\pi_{\theta}$ by maximizing the following clipped surrogate objective:
$\mathcal{J}_{\text{GRPO}}(\theta)=\;\mathbb{E}\left[\frac{1}{G}\sum_{i=1}^{G}(\min(r_{i}A_{i},\,\text{clip}(r_{i},1{-}\epsilon,1{+}\epsilon)A_{i})\right.\\ -\beta\,\mathcal{D}_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}}))\Bigg{]}$ (7)
where $r_{i}=\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)}$, $\epsilon$ controls clipping for stable optimization, and $\beta$ regulates the strength of the KL-divergence penalty with respect to a reference policy $\pi_{\text{ref}}$. Note that the expectation $\mathbb{E}[\cdot]$ is taken over the random variables $q$ and $\{o_{i}\}_{i=1}^{G}$, with $q\sim P(Q)$ and $\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O\mid q)$ and that the KL-divergence penalty encourages policy updates that remain close to the pretrained model:
$\mathcal{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})=\sum_{o_{i}}\pi_{\text{ref}}(o_{i}|q)\log\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-1.$ (8)
III-C Validation Module
The validation module ensures that injected code $c^{v}$ meaningfully introduces the intended vulnerability into the original benign function $c^{b}$. This process involves both deterministic tools and language model agents, and follows the workflow depicted in Figure 1.
First, a diff agent computes a token-level difference between $c^{v}$ and $c^{b}$ (Difference agent). If the difference is trivial (e.g., limited to whitespace, comments, or reformatting), the workflow returns to the injection stage with corrective feedback (Vuln. Injector Identical Code agent).
If the change is substantive, a vulnerability verification agent (LLM-based) assesses whether the injected code actually reflects the vulnerability that was intended (Diff. Analyzer agent). This agent is provided with rich context, including the original function $c^{b}$, the modified function $c^{v}$, their annotated difference, the vulnerability class to inject, and a self-rationale generated by the injection agent, which explains what edits were made and why they introduce the vulnerability. This enables the agent to evaluate not just the presence of code edits, but also whether those edits semantically fulfill the injection objective.
The modified code is analyzed using cppcheck [^1] https://cppcheck.sourceforge.io/. A validation agent then interprets its diagnostic report alongside prior context to assess whether the injected vulnerability aligns with the intended CWE (Vuln. Verifier).
If validation fails at this stage, the code is passed to a refinement agent, which revises the output based on prior feedback. The updated code is re-evaluated through the same pipeline. To ensure termination, the workflow is bounded by a fixed number of correction iterations: 10.
This hybrid validation strategy combines programmatic and rule-based tools, with LLM-based contextual reasoning, providing robust and interpretable guarantees over the correctness of injected vulnerabilities.
IV Experimental Setup
This section details the experimental setup designed to systematically evaluate the AVIATOR framework. We aim to answer the following research questions (RQs):
  • RQ1: How does AVIATOR perform compared to existing automated methods for creating realistic, large-scale vulnerability datasets?
  • RQ2: What impact does fine-tuning the vulnerability injection agent have on AVIATOR’s performance?
  • RQ3: How do different compositions of agents affect AVIATOR’s overall vulnerability generation performance?
IV-A Benchmark Construction
We evaluate AVIATOR specifically for C/C++ code, given the inherent susceptibility of these languages to low-level vulnerabilities (e.g., memory errors), focusing on the function level in alignment with prior vulnerability detection and repair studies [39]. We used three complementary popular open-source datasets used in previous studies [40] [32] [11].
  • SARD test suite 100 [16, 40]: Contains a set of 102 compact, synthetic test cases, including 34 vulnerable/benign function pairs that can be detected through formal verification tools such as ESBMC [33]. Due to the compact size and self-contained structure of these examples, we use SARD 100 to perform rapid fully automated validation.
  • FormAI [32]: Includes more realistic and structurally complex synthetic examples. While it does not provide aligned benign/vulnerable pairs, we manually created benign versions for 37 randomly sampled vulnerable functions, without altering the program’s intent and ensuring no vulnerabilities were detectable by formal verification tools on the resulting benign versions. ESBMC is again used for manual validation, offering a stronger challenge to our injection framework.
  • PrimeVul [11]: Contains over 5,000 real-world C/C++ code pairs annotated with vulnerability labels. Due to the lack of full context (e.g., missing class definitions, global variables, and interprocedural dependencies), PrimeVul is not suitable for automated formal verification. Instead, we use it for (1) fine-tuning the injection model, (2) constructing the retrieval-augmented generation (RAG) knowledge base, and (3) manually evaluating injection success on 45 randomly sampled pairs.
IV-B Implementation Details
Agents
AVIATOR leverages various technologies and tools. Within the validation workflow, we utilized the static analysis tool Cppcheck version 2.17, due to its popularity and ease of integration.
We experimented with two base large language models for the AI agents in our workflow. Qwen2.5-Coder-32B-Instruct [41], is a specialized open-source coding LLM with 32 billion parameters. We additionally experimented with Llama-4-Maverick [42], a general-purpose, mixture-of-experts-based LLM with approximately 400 billion parameters (17 billion active parameters), to explore the impact of model size and specialization on the agents’ reasoning capabilities.
Retrieval-Augmented Generation
We use gte-Qwen2-1.5B-Instruct for embedding vulnerable and benign function pairs in the RAG knowledge base. This embedder achieves top scores on the MTEB benchmark for embeddings of up to 8K tokens [43]. Since 99% of our training dataset from PrimeVul fits within this limit, we avoid chunking. Retrieval parameters were set at the recommended $k=4$.
Fine-Tuning Configurations
We fine-tune the vulnerability injection agent using 3.5k aligned pairs from the PrimeVul training set. For supervised fine-tuning (SFT), we trained for 5 epochs (loss convergence) with a learning rate of $2\times 10^{-4}$, batch size of 4, and LoRA ranks varying between 16 and 256. For reinforcement learning with GRPO, we trained the model for 1 epoch (reward convergence), using a learning rate of $2\times 10^{-4}$, batch size of 8, and generating four completions per policy iteration. The KL regularization coefficient $\beta$ (0.04) and clipping parameter $\epsilon$ (0.2) followed the recommended default values from GRPO implementation using Adam optimizer.
Technology
AVIATOR is implemented using Python as well as HuggingFace’s Transformers and Parameter-Efficient Fine-Tuning libraries. All experiments were run on NVIDIA A100 or H100 80GB GPUs.
IV-C Ablation Study Design
We conduct an ablation study to quantify the contribution of each agentic component in the AVIATOR workflow. To assess the necessity and impact of this modular structure, we construct and evaluate six progressively simplified variants by incrementally removing components from the full workflow.
The complete AVIATOR pipeline presented in Figure 1, showcases 13 distinct agents. These agents include tool-based nodes, LLM agents, and control logic designed to manage iterative processes.
The evaluated configurations are as follows:
  • 1-Agent (Monolithic LLM), $W_{1}$: A single LLM receives the original function, vulnerability type, and retrieved examples, and is prompted to analyze the function, inject the vulnerability, and self-verify. This setting relies on common chain-of-thought prompting without agentic decomposition.
  • 3-Agent, $W_{3}$: Decomposes the task into three distinct agents: (i) loading the vulnerability description, (ii) analyzing the function for injection points, and (iii) performing the injection. No verification or refinement stages are included.
  • 5-Agent (Injection-Only Module), $W_{5}$: Implements the full injection module (Code Understanding, Information Loader, Sanitization Detector, Vulnerability Injector, Polisher) without validation (Figure 1, top).
  • 7-Agent, $W_{7}$: Adds the difference-checking loop, consisting of the Diff Agent and Diff Analyzer, allowing the system to detect and iteratively refine the code if no substantive change was introduced.
  • 9-Agent, $W_{9}$: Adds the cppcheck static analysis tool and corresponding Verifier Agent to further validate the generated vulnerability and refine injection quality based on external tool feedback.
  • 11-Agent, $W_{11}$: Incorporates the Diff Refiner agent after the difference is computed, allowing targeted edits based on diff feedback. This step improves alignment between injection goals and actual edits by making minimal context-aware adjustments.
  • 13-Agent (Full AVIATOR), $W_{13}$: Adds a terminal Vulnerability Fixer Agent. This final pass ensures that injected vulnerabilities are realistic, contextually valid, and non-trivially removable—closely mimicking the post-injection review performed by security experts.
All configurations share the same injection model, prompt formatting, and RAG setup. We evaluate these variants on the same benchmark split (see Section 3, A) and report both formal and manual success metrics (see Section 3, D). This incremental comparison enables a fine-grained understanding of how individual agents affect the correctness, realism, and efficiency of the injection process.
IV-D Evaluation and Metrics
We assess the success rate of vulnerability injection, as it directly indicates the expected label accuracy of datasets produced by our framework
Automated Evaluation
For compatible test cases [16] and [32], we automatically measured vulnerability injection success using the formal verification tool ESBMC v7.5 (Efficient SMT-based Bounded Model Checker), recognized for its effectiveness in detecting vulnerabilities in bounded code contexts [33, 44]. It implements state-of-the-art incremental Bounded Model Checking and $k$-induction proof-rule algorithms. Namely, if the solver reaches termination within a bound less than or equal to $k$, we can definitively prove the absence of software errors. We set the bound to 100 and timeout to 5 minutes.
Due to the probabilistic nature of LLMs used in AVIATOR, identical inputs may yield different outputs on repeated trials. Therefore, we generated 5 outputs for SARD100 and FormAI test cases. To rigorously evaluate AVIATOR’s performance, we employed two metrics:
  • Average Injection Success Rate (AISR): Measures the proportion of generated vulnerabilities correctly injected and verified, averaged across multiple runs. Denote $s_{i}$ as the success rate in run $i$ over $n$ runs:
    $AISR_{n}=\frac{1}{n}\sum_{i=1}^{n}s_{i},\quad\sigma=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(s_{i}-\text{AISR})^{2}},$ (9)
    where $\sigma$ denotes the empirical standard deviation.
  • Pass@k [45]: Estimates the probability that at least one of $k$ independently sampled outputs is correct. If $c$ is the number of correct outputs among $k$ samples:
    $\text{Pass@}k=1-\frac{\binom{n-c}{k}}{\binom{n}{k}},\quad\text{if }c<k;\quad\text{Pass@}k=1\text{ otherwise}$
    This metric captures the practical success likelihood under multi-sample inference.
These metrics provide insights into AVIATOR’s reliability and practical utility, reflecting its robustness and stability in automated vulnerability injection tasks.
Manual Evaluation
For PrimeVul, we manually validated a sample of 45 injected functions drawn uniformly at random, following prior work [21, 10, 29, 46]. The evaluation was performed by three authors experienced in C/C++, including two senior security researchers. Each injected sample was categorized as (1) weakness present, (2) no weakness, or (3) undecidable—typically due to reliance on external functions or context beyond the injected function. Since we lack complete test cases for most programs, we assess the presence of weaknesses rather than confirmed vulnerabilities (which require demonstrable exploitability). The reported injection success rate reflects the percentage of samples found to contain a weakness, based only on those that could be conclusively assessed.
V Results and Analysis
In the following we investigate the answers to the research questions posed earlier.
V-A RQ1. Effectiveness of AVIATOR Compared to Existing Dataset Creation Techniques
To evaluate AVIATOR’s effectiveness, we assessed its vulnerability injection success using samples from three benchmarks: SARD100, FormAI, and PrimeVul. Table I shows the average success rates for the full AVIATOR workflow ($W_{13}$) with the SFT injection agent. AVIATOR achieved a 95% success rate on SARD100 and 91% on the more complex FormAI benchmark, indicating strong generalization beyond simple patterns.
For PrimeVul, manual analysis was conducted on 45 injected samples due to the lack of external context needed for automated verification. Out of 34 analyzable cases, 32 contained a verifiable weakness within the function body (94%). The remaining samples involved undefined external symbols, so the analysis focused on confirming the presence of weakness insertion rather than evaluating full vulnerability exploitability.
TABLE I: AVIATOR injection success across datasets: Average Injection Success Rate on 5 runs for SARD100 and FormAI, and on analyzable cases for PrimeVul due to limited context.
Dataset Eval. Type Metric AVIATOR
SARD100 Automated $AISR_{5}$ 95%
FormAI Automated $AISR_{5}$ 89%
PrimeVul Manual % Analyzable Weaknesses 94%
Table II compares AVIATOR with other techniques, showing that its success rate on FormAI (91%) and PrimeVul (94%) surpasses existing injection methods, like VinJ and VulGen, which report 69% success rates. AVIATOR’s guided injection process allows for precise, context-aware edits that mimic developer like mistakes, making it effective even in partially defined or real-world settings.
TABLE II: Label accuracy and injection success rates (marked with *) for major vulnerability datasets that adhere to BCD properties (see Section II-A). AVIATOR results are based on automated assessment using the FormAI dataset and manual validation using PrimeVul.
Model / Dataset (*) Accuracy Source
Graph2Edit * [47] 13% [28]
BigVul [17] 25% [11]
AVIATOR $W_{1}$ * 31% FormAI validation
CrossVul [19] 48% [11]
Getafix * [47] 50% [28]
CVEFixes [18] 52% [11]
DiverseVul [10] 60% [11]
VulGen * [28] 69% [28]
VinJ * [29] 69% [29]
AVIATOR $W_{13}$ & SFT * 91% FormAI validation
AVIATOR $W_{13}$ & SFT * 94% PrimeVul validation
Answer to RQ1: AVIATOR achieves high injection success across benchmarks: 95% on Sard100, 91% on FormAI, and 94% verifiable weakness insertion on PrimeVul. These results highlight AVIATOR’s strong generalization and realism, even in partially specified or real-world code settings. We attribute this performance to AVIATOR’s novel use of an agentic workflow that incorporates multi-step guiding and task decomposition, that enables the injection agent to iteratively refine and validate its own output, likely contributing to its effectiveness across diverse benchmarks.
V-B RQ2. Impact of Fine-Tuning the Injection Agent on AVIATOR’s Performance
To assess the effect of model-level adaptation, we evaluate AVIATOR’s injection performance before and after fine-tuning the injection agent. We experiment with two fine-tuning regimes: supervised fine-tuning (SFT) and reinforcement learning via Grouped Relative Policy Optimization (GRPO), using the same base model (Qwen2.5-Coder-32B) and the full agentic pipeline ($W_{13}$).
Crucially, all fine-tuning is performed exclusively on PrimeVul—a dataset that shares no overlap with either SARD100 or FormAI. This ensures the evaluation measures true generalization, not memorization.
SFT is both more effective and more efficient than GRPO
Table III reports the average injection success rate ($AISR_{5}$) and its standard deviation across 5 vulnerability injection runs on the same data. On the more complex FormAI dataset, SFT yields significant gains over the base model (91% vs. 85%), while also reducing variance. In contrast, GRPO not only fails to outperform the base model on FormAI (dropping to 84%), but also provides no gain on SARD100. These results indicate that in our setup, GRPO is less effective than SFT, despite using the same training data and more training computation.
On the simpler SARD100 benchmark, the gains from SFT are marginal (95% vs. 94%), indicating that AVIATOR’s base configuration already performs near-saturation. However, the slight improvement and reduced variance still confirm the benefit of tuning.
TABLE III: AVIATOR $W_{13}$ injection success rate ($AISR_{5}$) across 5 vulnerability injection generation with and without fine-tuning, on FormAI and SARD100.
Dataset Model Variant $AISR_{5}$ Std
FormAI Qwen2.5 (no FT) 85% 4
Qwen2.5 + GRPO 84% 1
Qwen2.5 + SFT 91% 1
SARD100 Qwen2.5 (no FT) 94% 3
Qwen2.5 + GRPO 94% 0
Qwen2.5 + SFT 95% 2
Pass@k Improvements with SFT
We also reported Pass@k performance to assess the likelihood of producing a successful injection within $k$ generation attempts. Table IV shows that SFT improves results consistently across all $k$ on both benchmarks. On FormAI, Pass@1 improves by over 5 percentage points (from 84.3% to 89.9%), and SARD100 reaches near-perfect performance even for small $k$. These gains further validate the robustness added by SFT.
TABLE IV: AVIATOR $W_{13}$ pass@k for 10 vulnerability injection generations with and without fine-tuning, on FormAI and SARD100.
Dataset Model @1 @2 @3 @4 @5 @6 @7 @8 @9 @10
FormAI $W_{13}$ 84.32% 91.05% 93.87% 95.24% 96.03% 96.56% 96.94% 97.18% 97.30% 97.30%
$W_{13}$ + SFT 89.91% 93.04% 94.81% 95.62% 96.15% 96.55% 96.87% 97.14% 97.38% 97.58%
SARD100 $W_{13}$ 93.58% 95.95% 96.66% 96.91% 97.01% 97.05% 97.06% 97.06% 97.06% 97.06%
$W_{13}$ + SFT 95.15% 98.65% 99.51% 99.79% 99.93% 99.99% 100.00% 100.00% 100.00% 100.00%
Answer to RQ2. Fine-tuning the vulnerability injection agent using SFT leads to consistent and significant performance gains for the whole AVIATOR workflow, particularly on complex datasets like FormAI. GRPO fails to match this improvement despite being substantially more resource-intensive. These results highlight the efficiency and generalization strength of SFT, even when trained on an entirely separate dataset.
V-C RQ3. Contribution of Agentic Components and Model Variants to Injection Performance
To evaluate the contribution of AVIATOR’s agentic workflow design, we performed an ablation study on seven progressively extended configurations ($W_{1}$$W_{13}$), incrementally adding specialized agents for analysis, validation, and refinement. We additionally compare four model variants: (i) base Qwen2.5-Coder, (ii) Qwen2.5 fine-tuned using SFT, (iii) Qwen2.5 fine-tuned using GRPO, and (iv) the general-purpose Llama-4-Maverick. Figure 2 shows the average injection success rate across 5 trials ($AISR_{5}$) on both SARD100 and FormAI.
Performance Improves with Workflow Depth
Across all models, performance increases consistently as more specialized agents are added. For example, with the base Qwen2.5-Coder, success rises from 31% at $W_{1}$ to 85% at $W_{13}$ on FormAI, and from 84% to 94% on SARD100. The benefits of modular reasoning are particularly pronounced on complex tasks (FormAI), confirming that decomposing the task into subtasks handled by agents enhances robustness and clarity.
Refer to caption Refer to caption
(a) SARD 100 (b) FormAI
Figure 2: AVIATOR average injection success rate on 5 attempts. Ablation study for LLM agents using Llama-4-Maverick, Qwen2.5-Coder without fine-tuning and Qwen2.5-Coder with the vulnerability injection agent fine-tuned using SFT and GRPO.
Impact of Specific Agentic Additions
Moving from $W_{1}\rightarrow W_{3}$ introduces basic task decomposition, yielding large but unstable gains on FormAI (+40%) due to better prompt structuring. At $W_{5}$, AVIATOR includes the complete injection pipeline, stabilizing outputs and significantly improving accuracy (+5–10%). Furthermore, by adding feedback-based validation via diff checking ($W_{7}$) and static analysis with cppcheck ($W_{9}$) further boosts performance, particularly for models with more specialized abilities like the fine-tuned Qwen2.5-Coder. Final refiners ($W_{11}$, $W_{13}$) provide consistent final gains across models, enabling subtle bug integration and polishing.
SFT and GRPO Validated Across Configurations
To ensure consistent comparisons, we fine-tuned the vulnerability injection agent starting from $W_{5}$, where the agent’s input structure and prompt format remain unchanged across workflows. Configurations below $W_{5}$ ($W_{1}$, $W_{3}$) are not compatible with our fine-tuned models, as they would require new training due to different agent prompts and inputs. Consistent with results in RQ2, SFT delivers the highest gains across all configurations, while GRPO fails to improve over the base model despite significantly longer training time.
General-Purpose vs. Specialized Models
Llama-4-Maverick, a 400B-parameter generalist model, underperforms all Qwen2.5-Coder 32B variants across the board. For example, on FormAI at $W_{13}$, it reaches 77% compared to 85% with the base Qwen2.5 and 91% with SFT. This underlines the advantage of coding-specialized models and domain-specific fine-tuning for structured tasks like vulnerability injection.
Answer to RQ3. Injection performance improves with workflow refinement: from 31% at $W_{1}$ to 91% at $W_{13}$ on FormAI. Fine-tuning the injection agent using SFT amplifies these gains even though most of the base performance comes from a well-crafted workflow. GRPO fails to outperform the base model. Domain-specialized models (e.g., Qwen2.5-Coder) clearly outperform general-purpose LLMs (e.g., Llama-4), demonstrating the value of both specialization and expert-based agentic decomposition.
VI Discussion
This section reflects on why AVIATOR works, its practical impact, limitations, and directions for future research.
VI-A Why Does AVIATOR Work?
We identify the core design choices that explain AVIATOR’s effectiveness in vulnerability injection. These include its agentic decomposition, targeted model-level adaptation, iterative refinement capabilities, and integration of static analysis feedback.
(1) AVIATOR benefits from agentic decomposition. The injection task is divided among 13 specialized agents, each handling specific operations like vulnerability localization, injection, sanitization detection, or diff-based verification. This modular approach enhances accuracy significantly, as demonstrated by improvements from 31% to 91% on FormAI and from 84% to 94% on SARD100.
(2) AVIATOR improves through targeted fine-tuning of a single critical agent. Rather than retraining all models, we fine-tune only the Vulnerability Injection agent using supervised fine-tuning (SFT) with LoRA. This lightweight adaptation requires less than 10 hours of training on a single NVIDIA A100 GPU and improves injection accuracy from 85% to 91% on FormAI. The model is trained on PrimeVul, a dataset entirely disjoint from the test sets, showing that this adaptation generalizes well across domains.
(3) AVIATOR leverages iterative workflow refinement to improve injection robustness. Agents operate in feedback loops to enhance injection robustness. For example, the Diff Agent triggers refinement cycles when necessary, boosting accuracy up to 10 points on FormAI and 4 points on SARD100.
(4) AVIATOR incorporates traditional static analysis tools to validate injection quality. The Verifier Agent uses cppcheck to detect whether the injected code violates real-world safety constraints. Its feedback is interpreted and used by subsequent agents to iteratively refine the injection.
VI-B Limitations and Threats to Validity
Benchmark Scope and Generalizability
Our evaluation focuses on C/C++ using three datasets (SARD100, FormAI, and PrimeVul), which may not capture all real-world diversity. To mitigate this, we combined synthetic and real-world benchmarks of varying complexity.
Manual Evaluation Bias
The qualitative analysis of PrimeVul is based on 45 manually annotated samples, which may introduce subjectivity. We mitigated this by involving three security researchers and resolving disagreements via consensus.
Metric Stability Under Sampling
LLM generation is stochastic, which may affect reliability of metrics like injection success rate. To address this, we performed multiple independent runs and reported average scores $AISR_{k}$ with standard deviations across all experiments and measured Pass@k with additional runs on the best configurations.
Generalization of Fine-Tuned Agent
Our supervised fine-tuning was performed solely on PrimeVul, raising a potential risk of overfitting to its code characteristics. However, PrimeVul includes a wide variety of real-world vulnerabilities across multiple domains. Moreover, the consistent performance improvements on structurally distinct benchmarks (SARD100 and FormAI) provided strong evidence of generalization beyond the training distribution.
VII CONCLUSION
In this paper, we introduced AVIATOR, an AI agentic framework designed to systematically inject realistic and category-specific vulnerabilities into secure codebases. By leveraging a multi-agent pipeline that replicates the reasoning processes of cybersecurity experts, AVIATOR ensures that vulnerability injection is not only precise but also deeply contextually aware. This approach allows for a nuanced understanding and handling of vulnerabilities, closely mimicking the analytical depth and precision of human experts.
Our experimental evaluations, conducted across diverse benchmarks, demonstrate that AVIATOR seems to significantly outperform existing techniques, achieving high success rates in vulnerability injection. The modular design of AVIATOR, combined with cutting-edge techniques such as Retrieval-Augmented Generation and Low-Rank Adaptation, enables the creation of large-scale datasets. These datasets are not only accurately labeled but are diverse, making them very useful for training robust and effective vulnerability detection models.
The contributions of this work not only advance the state-of-the-art in automated vulnerability injection but also provide a scalable solution to the pressing challenge of data scarcity in software security research.
Future work will explore further enhancements in agentic workflows and fine-tuning strategies, as well as further evaluation on real-world code, to broaden the scope and applicability of AVIATOR in practical scenarios.